if x+2y=10andx+5y=19what is the value of y
Answer by edjones(7311)

A) x+2y=10
B) x+5y=19
.
A) x=10-2y subtract 2y from each side.
B)10-2y+5y=19
3y=9 add -10 to each side.
y=3 divide 3 by each side.
.
Ed